# Towards Calibrated Robust Fine-Tuning of Vision-Language Models

### v5

@@ Abstract
Improving out-of-distribution (OOD) generalization through in-distribution (ID) adaptation is a primary goal of robust fine-tuning methods beyond the naive fine-tuning approach. However, despite decent OOD generalization performance from recent robust fine-tuning methods, OOD confidence calibration for reliable machine learning has not been fully addressed. This work proposes a robust fine-tuning method that improves both OOD accuracy and calibration error in Vision Language Models (VLMs). Firstly, we show that both types of errors have a shared upper bound consisting of two terms of ID data: 1) calibration error and 2) the smallest singular value of the input covariance matrix. Based on this insight, we design a novel framework that conducts fine-tuning with a constrained multimodal contrastive loss enforcing a larger smallest singular value, which is further aided by the self-distillation of a moving averaged model to achieve well-calibrated prediction. Starting from an empirical validation of our theoretical statements, we provide extensive experimental results on ImageNet distribution shift benchmarks that demonstrate the effectiveness of our method. Our code is available here.
**footnotetext: Equal contribution (changdae.oh@uos.ac.kr; hyesulim@kaist.ac.kr)++footnotetext: Co-correspondence (zhiqic@cs.cmu.edu; kyungwoo.song@yonsei.ac.kr)
@@ 1 Introduction
Foundation models [5] have been extensively used to fine-tune downstream tasks using a pre-training-fine-tuning approach, employing generalized knowledge from large pre-training data to achieve both in-distribution (ID) and out-of-distribution (OOD) performance. After Wortsman et al. [53] and Kumar et al. [25] amplified the discussion on the robustness of recent Vision-Language Models (VLM)s [40, 22, 39] under distribution shifts, a wide range of research has followed [26, 13, 26, 49, 37] so-called robust fine-tuning.
Figure 1: OOD accuracy vs. ID accuracy (left) and negative OOD ECE (right). Notice that the competing methods – FLYP [13], LP-FT [25], and Lipsum-FT [36] – improve generalization over the zero-shot baseline (ZS) and naive fine-tuning (FT) but suffer from OOD miscalibration, presumably due to concerning model generalization solely at fine-tuning. Our CaRot outperforms existing methods on both OOD generalization and calibration by large margins. To maintain consistency in the plots, where desired values are shown on the right side of the x-axis, we report negative OOD ECE. ID ACC refers to ImageNet-1K top-1 accuracy; OOD ACC and ECE refer to the averaged accuracy and ECE of the five ImageNet distribution shifts (ImageNetV2, ImageNet-R, ImageNet-A, ImageNet-Sketch, and ObjectNet), respectively. Detailed numbers are reported in Table 2 and 3.
Despite the advancements of the methods, we find that an important criterion for trustworthy machine learning (ML) has been overlooked – confidence calibration [33, 7] that quantifies how close the confidence of our predictor is to the actual correctness of predictions. As shown in Figure 1, existing robust fine-tuning methods hurt the confidence calibration in terms of expected calibration error (ECE) [34] on OOD data while they show improvements on OOD accuracy compared to the zero-shot performance. In this work, we introduce a calibrated robust fine-tuning method (CaRot) that simultaneously improves calibration and accuracy on OOD data.
@@
To accomplish low classification and calibration errors on OOD samples when only ID samples are available, we conduct theoretical analyses of those OOD errors. To be specific, we derive an upper bound that is shared with OOD classification error and OOD calibration error composed with two quantities on ID samples: 1) the reciprocal of the smallest singular value of the normalized covariance matrix of data representation and 2) the calibration error. Different from the existing bounds focusing on either one of classification or calibration error [4, 63, 54], we address both classification and calibration errors in a single unified bound. It is worth noting that we can compute the chief components of our bound without relying on unlabeled OOD samples [3].
@@
Motivated by these theoretical analyses, we propose a new multimodal contrastive loss that promotes the smallest singular value of input image representation to become larger by enforcing the orthogonality of the final linear projection layer. Furthermore, to understand the proposal’s working mechanism in depth, we present an interpretation of our new multimodal contrastive loss as a process of seeking the low-rank approximation of cross-covariance matrix over image-text representations with reduced parameter space via orthogonality constraint. Meanwhile, to enhance confidence calibration on ID samples during fine-tuning, we utilize a self-distillation (SD) with an exponential moving average (EMA) teacher model. This EMA SD encourages a student model to learn semantic similarity structures of representations from teacher predictions across diverse contrastive pairs, appropriately adjusting confidence per each instance.
@@
We first validate our new error bounds with synthetic datasets to show that the bounds hold empirically. Then, evaluate our method by conducting extensive experiments of fine-tuning CLIP [40] on ImageNet-1K [8] classification task under natural shift (ImageNet-V2/R/A/Sketch and ObjectNet). We demonstrate the effectiveness of our proposed framework for robust generalization and calibration by observing significant improvement in terms of expected calibration error and accuracy on ID/OOD datasets.
@@
Contributions. 1) We point out that existing fine-tuning methods do not adequately achieve satisfactory OOD generalization and calibration simultaneously. 2) We provide theoretical analysis for classification and calibration errors on the OOD domain and show that they are both bounded from above by ID calibration error and the smallest singular value of the covariance matrix over ID input representation. 3) Based on our theoretical analyses, we devise a calibrated robust fine-tuning method, CaRot, that reduces the upper bound of OOD classification and calibration errors by conducting constrained multimodal contrastive learning with EMA self-distillation. 4) We first present empirical evidence for our theory on a synthetic dataset and demonstrate the efficacy of CaRot via extensive evaluations on ImageNet-1K distribution shifts in terms of accuracy and calibration error on ID and OOD domains.
@@ 2 Preliminary
Robust fine-tuning aims to achieve consistently high performance on both training (ID) and related but different test distributions (OOD). For validation, we commonly consider a covariate shift scenario for the classification task, where both ID and OOD domains share the class categories ($\mathcal{Y}_{\text{ID}}=\mathcal{Y}_{\text{OOD}}$) and have the same conditional distribution $P(Y|X)$, but have different marginal distributions over input $X$. That is, $P_{\text{ID}}(X)\neq P_{\text{OOD}}(X)$ but $P_{\text{ID}}(Y|X)=P_{\text{OOD}}(Y|X)$. Here, we evaluate the robustness of a model, which is fine-tuned on a training split of ID domain, on a test split of ID and that of OOD domains.
Confidence calibration is a concept of matching the softmax probabilities outputted for different inputs to the expected accuracy on these inputs. In a $K$-way classification setting, let $X\in\mathbb{R}^{d}$ and $Y\in\{1,...,K\}$ be random variables indicating inputs and labels, respectively. A dataset with $N$ independent samples from the joint distribution $P(X,Y)=P(Y|X)P(X)$ is denoted by $\{(x_{n},y_{n})\}^{N}_{n=1}$. Let $f$ be a classifier and $f(y|x)=\hat{p}$ be a confidence, i.e., the maximum of output probability corresponding to its prediction $\hat{y}$. We say a model is perfectly-calibrated when $\mathbb{P}(\hat{y}=y|\hat{p}=p)=p,\ \forall p\in[0,1]$. To measure the calibration of a model, the model calibration can be derived as $\mathbb{E}[|\mathbb{P}(\hat{y}=y|\hat{p}=p)-p|]$. In practice, we use expected calibration error (ECE) [34] as an empirical approximation of the model calibration, which is a weighted average of bin-wise miscalibration. The ECE divides the confidence score of $N$ samples into $M$ uniform confidence bins $\{B_{m}\}_{m=1}^{M}$ and takes the mean of the gap between accuracy (acc) and confidence (conf) over the bins weighted by the number of samples in the bins, i.e.,$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|\text{acc}(B_{m})-\text{conf}(B_{m})|$.
@@ 3 Theoretical Analysis on OOD Generalization and Calibration
We first identify the factors that affect OOD generalization and calibration errors under circumstances where only ID data is accessible. We take inspiration from the generalization bound of domain adaptation literature [4, 63] while remarkably adapting the analysis to consider both OOD classification and calibration error at the same time in a more practical way.
Let $\mathcal{D}$ be a domain on input space $\mathcal{X}$ and $\mathcal{Y}=\{0,1\}$ be a label space for a binary classification. Among the hypothesis function $h:\mathcal{X}\rightarrow[0,1]$, we define $h_{0}(\cdot)$ as a desired calibrated predictor for label $y$, which minimizes the calibration error $\mathbb{E}_{x\sim\mathcal{D}}[(h(x)-c(x))^{2}]$ [33], where $c(x)=\mathbb{E}_{y}[y|h(x)]$ is the expected value of $y$ given a prediction $h(x)$. That is, $h_{0}$ always produces the calibrated prediction for $y$ given $x$ so that the output confidence $h_{0}(x)$ matches the expectation of $y$ over the subset of samples that have the same confidence as $h_{0}(x)$. Our goal is to learn a hypothesis function $h(\cdot)$ that outputs reliable probability on samples from the unseen OOD domain, which is defined by a distribution $\mathcal{D}_{\text{OOD}}$, as well as on the ID domain $\mathcal{D}_{\text{ID}}$, where the predictor is trained on. In essence, the error $\varepsilon_{\mathcal{D}}(h)=\mathbb{E}_{x\sim\mathcal{D}}[(h(x)-h_{0}(x))^{2}]$ should be small for two different domains $\mathcal{D}\in\{\mathcal{D}_{\text{ID}},\mathcal{D}_{\text{OOD}}\}$. In this work, we concentrate on the covariate shift scenario (§2) that only the marginal distribution over $\mathcal{X}$ changes while the distribution over $\mathcal{Y}$ is preserved.
Let the optimal hypothesis $h^{*}$, which minimizes a combination of errors on both ID and OOD domain, be $h^{*}:=\operatorname*{arg\,min}_{h\in\mathcal{H}}\varepsilon_{\mathcal{D}_{\text{ID}}}(h)+\varepsilon_{\mathcal{D}_{\text{OOD}}}(h)$, and $\Delta$ denote the optimal joint error $\Delta:=\varepsilon_{\mathcal{D}_{\text{ID}}}(h^{*})+\varepsilon_{\mathcal{D}_{\text{OOD}}}(h^{*})$. We derive our novel bounds for the OOD calibration error and OOD classification error.
@@
Let $h:\mathcal{X}\rightarrow[0,1]$ be a real-valued function which has structure $h(x)=\sum_{i=1}^{d}h_{i}(x_{i})$ where $h_{i}$ is an arbitrary one-dimensional function, and $h(\cdot)$ is in a hypothesis class $\mathcal{H}$ that has pseudo dimension $\mathcal{P}dim(\mathcal{H})=d_{h}$, $\hat{\mathcal{D}}_{\text{ID}}$ be an $N$-size empirical distribution on ID domain. If $(x_{i},x_{j})$ is a bi-variate Gaussian random variable for every $i,j\in[d]$, then for any $\delta\in(0,1)$ and for all $h$, the following bounds hold with probability at least $1-\delta$:
@@
where $\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}}{:=}\mathbb{E}_{\mathcal{D}_{\text{ID}}}[\tilde{x}\tilde{x}^{T}]$ is a covariance matrix of $d$-dimensional normalized input $\tilde{x}=(\tilde{x}_{1},...,\tilde{x}_{d})$, where $\tilde{x}_{i}{:=}(x_{i}-\mathbb{E}[x_{i}])\text{Var}(x_{i})^{-1/2}$ and $\sigma_{\text{min}}(M)$ is the smallest singular value of a matrix $M\in\mathbb{R}^{d_{1}\times d_{2}}$. These theoretical results can be directly applied to the intermediate representation of a neural network by setting the input variable $x$ as a representation vector [3, 64], so we will assume the input as an image representation of the visual encoder in following sections. Note that 1) the LHS of the first inequality (eq.(1)) indicates OOD calibration error, 2) two terms in the LHS of the second inequality (eq.(2)) denote OOD classification error in terms of $L_{2}$ loss and prediction sharpness in OOD domain, and 3) both inequalities have the same RHS, which contains the empirical estimate of ID calibration error, the reciprocal of the smallest singular value of ID input covariance matrix, and remaining irreducible terms that depend on the problem setup.
@@ 4 Method
Figure 2: Overview of CaRot. We fine-tune VLM using a multimodal contrastive loss with an orthogonality constraint on visual projection layer (eq.(4)) and self-distillation $\mathcal{L}_{\text{SD}}$ (eq.(5)) that takes predictions of EMA teacher $\psi$ as soft target labels to update the student model $\theta$. The darker and the lighter elements denote values closer to 1 and 0, respectively. Both teacher and student models share identical VLM architecture consisting of image $f_{\theta_{v}}:=[f_{\hat{\theta}_{v}};W_{v}]$ and text $g_{\theta_{l}}:=[g_{\hat{\theta}_{l}};W_{l}]$ encoders, where $W$ is the last projection layer. Given (image, text) pair data, the model outputs the similarity map across image-text representations.
Our goal is to achieve OOD generalization and calibration through ID adaptation of pre-trained VLMs. Motivated by our analysis in §3, we propose a calibrated robust fine-tuning method that enlarges the smallest singular value of the covariance matrix over ID input representation while improving ID calibration, thereby lowering the upper bound of OOD calibration and generalization error. Following previous works [53, 25, 13], we set CLIP [40] as our target VLM and focus the scope of validation to image classification task. An overview of our method is illustrated in Figure 2.

### v7

@@ Abstract
Improving out-of-distribution (OOD) generalization during in-distribution (ID) adaptation is a primary goal of robust fine-tuning of zero-shot models beyond naive fine-tuning. However, despite decent OOD generalization performance from recent robust fine-tuning methods, confidence calibration for reliable model output has not been fully addressed. This work proposes a robust fine-tuning method that improves both OOD accuracy and confidence calibration simultaneously in vision language models. Firstly, we show that both OOD classification and OOD calibration errors have a shared upper bound consisting of two terms of ID data: 1) ID calibration error and 2) the smallest singular value of the ID input covariance matrix. Based on this insight, we design a novel framework that conducts fine-tuning with a constrained multimodal contrastive loss enforcing a larger smallest singular value, which is further guided by the self-distillation of a moving-averaged model to achieve calibrated prediction as well. Starting from empirical evidence supporting our theoretical statements, we provide extensive experimental results on ImageNet distribution shift benchmarks that demonstrate the effectiveness of our theorem and its practical implementation. Our code is available here.
[…]
@@ 1 Introduction
Foundation models [6] such as CLIP [47] have been extensively utilized on diverse domains via pretrain-finetune approaches. Their generalized knowledge shaped after large-scale pre-training enables them to easily adapt to downstream tasks through zero-shot inference or fine-tuning. However, it has been steadily reported that a naive fine-tuning approach comprises foundation models’ strong out-of-distribution (OOD) generalization capability during adaptation to in-distribution (ID) data [61, 30]. To ensure robustness under distribution shifts, a wide range of research has followed [61, 30, 17, 32, 57, 43, 42] so-called robust fine-tuning.
Figure 1: OOD accuracy vs. ID accuracy (left) and negative OOD ECE (right). To maintain consistency in the plots, where desired values are shown on the right side of the x-axis, we report negative OOD ECE. ID ACC refers to ImageNet-1K top-1 accuracy; OOD ACC and ECE refer to the averaged accuracy and ECE of the five ImageNet distribution shifts (ImageNetV2, ImageNet-R, ImageNet-A, ImageNet-Sketch, and ObjectNet), respectively. Detailed numbers are reported in Table 2 and 3. Note that the competing methods – FLYP [17], LP-FT [30], and Lipsum-FT [42] – improve OOD accuracy over the zero-shot baseline (ZS) and naive fine-tuning (FT) but suffer from OOD miscalibration, presumably due to concerning generalization solely during fine-tuning. Our CaRot outperforms existing methods on both OOD accuracy and calibration by large margins.
Despite the advancements of the robust fine-tuning methods, we are aware that an important criterion for trustworthy machine learning has been overlooked – confidence calibration [39, 8] that quantifies how close the confidence of our predictor is to the actual correctness of predictions. As shown in Figure 1, existing robust fine-tuning methods hurt the confidence calibration in terms of expected calibration error (ECE) [40] on OOD data compared to the zero-shot evaluation while they show improvements on OOD accuracy. In this work, we introduce a calibrated robust fine-tuning method (CaRot) that simultaneously improves confidence calibration and accuracy of the classifier on OOD data.
@@
To accomplish low classification and calibration errors on OOD samples with only ID samples in our hands, we conduct theoretical analyses of those OOD errors. To be specific, we derive an upper bound that is shared for OOD classification error and OOD calibration error composed with two quantities on ID samples, 1) the reciprocal of the smallest singular value of the normalized covariance matrix of ID data representation and 2) the ID calibration error. Different from the existing bounds focusing on either one of classification or calibration error [5, 72, 63], we address both classification and calibration errors in a single unified bound. More importantly, the chief components of our bound can be computed solely with ID samples without relying on any OOD samples, which discerns our approach to existing work [4].
@@
Motivated by our theoretical analysis, we propose a new multimodal contrastive loss that promotes the smallest singular value of input image representation to become larger by enforcing the orthogonality of the final projection matrix of the visual encoder. Furthermore, to understand the working mechanism in depth, we present an interpretation of our new multimodal contrastive loss as a process of seeking the low-rank approximation of cross-covariance matrix over image-text representations on a reduced solution space induced by the orthogonality constraint. Meanwhile, to enhance confidence calibration on ID samples during fine-tuning, we utilize a self-distillation (SD) with an exponential moving average (EMA) teacher model. This EMA SD encourages a student model to learn semantic similarity structures of in-batch data representations from teacher predictions across diverse contrastive pairs, appropriately adjusting confidence per instance.
@@
We first validate our new error bounds with synthetic datasets to show that the bounds hold empirically. Then, evaluate our method by conducting extensive experiments of fine-tuning CLIP [47] on ImageNet-1K [10] classification task under natural distribution shift (ImageNet-V2/R/A/Sketch and ObjectNet) and synthetic distribution shift (ImageNet-C). We demonstrate the effectiveness of our proposed framework for robust generalization and calibration by observing consistent improvements in terms of expected calibration error and accuracy on ID/OOD datasets.
@@
Summary of contributions. 1) We point out that existing fine-tuning methods do not adequately achieve satisfactory OOD generalization and calibration simultaneously. 2) We provide theoretical analysis for classification and calibration errors on the OOD data and show that they are both bounded from above by the ID calibration error and the smallest singular value of the covariance matrix over the ID input representation. 3) Based on our theoretical analyses, we devise a calibrated robust fine-tuning method, CaRot, as a practical realization of our theorem that reduces the upper bound of OOD classification and calibration errors by conducting constrained multimodal contrastive learning with EMA self-distillation. 4) We present empirical evidence for our theory on a synthetic dataset and demonstrate the efficacy of CaRot via extensive evaluations on ImageNet-1K distribution shifts in terms of accuracy and calibration error on ID and OOD domains.
@@ 2 Preliminary
Robust fine-tuning aims to achieve consistently high performance on data from both training distribution (ID) and related but different test distributions (OOD). For validation, we commonly consider a covariate shift scenario for the classification task, where both ID and OOD domains share the class categories ($\mathcal{Y}_{\text{ID}}=\mathcal{Y}_{\text{OOD}}$) and have the same conditional distribution $P(Y|X)$, but have different marginal distributions over input $X$. That is, $P_{\text{ID}}(Y|X)=P_{\text{OOD}}(Y|X)$ but $P_{\text{ID}}(X)\neq P_{\text{OOD}}(X)$. Here, we evaluate a model that is fine-tuned on a training split of the ID domain, on a test split of ID, and on OOD domains. The term “OOD" is quite general, and we confine the scope of OOD to transformed and related distributions with ID [12]. For example, if our ID data is about an object recognition task with images, OOD data is about a sensor-noised version of ID, or independently collected data from a different domain targeting the same task.
Confidence calibration is a concept of matching the prediction probabilities yielded for different inputs to the expected accuracy on these inputs. In a $K$-way classification setting, let $X\in\mathbb{R}^{d}$ and $Y\in\{1,...,K\}$ be random variables indicating inputs and labels, respectively. A dataset with $N$ independent samples from the joint distribution $P(X,Y)=P(Y|X)P(X)$ is denoted by $\{(x_{n},y_{n})\}^{N}_{n=1}$. Let $f$ be a classifier and $f(y|x)=\hat{p}$ be a confidence, i.e., the maximum of probabilities among $K$ dimensions corresponding to its prediction $\hat{y}$. We say a model is perfectly-calibrated when $\mathbb{P}(\hat{y}=y|\hat{p}=p)=p,\ \forall p\in[0,1]$. As a quantitative measure, the model calibration can be derived as $\mathbb{E}[|\mathbb{P}(\hat{y}=y|\hat{p}=p)-p|]$. In practice, we use expected calibration error (ECE) [40] as an empirical approximation of the model calibration, which is a weighted average of bin-wise miscalibration. The ECE divides the confidence score of $N$ samples into $M$ uniform confidence bins $\{B_{m}\}_{m=1}^{M}$ and takes the mean of the gap between accuracy (acc) and confidence (conf) over the bins weighted by the number of samples in the bins, i.e., $\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|\text{acc}(B_{m})-\text{conf}(B_{m})|$.
@@ 3 Theoretical Analysis on OOD Generalization and Calibration
We first identify the factors that affect OOD generalization and calibration errors under circumstances where only ID data is accessible. We take inspiration from the generalization bound of domain adaptation literature [5, 72] while remarkably adapting the analysis to consider both OOD classification error and OOD calibration error at the same time in a more practical way.
Let $\mathcal{D}$ be a domain on input space $\mathcal{X}$ and $\mathcal{Y}=\{0,1\}$ be a label space for a binary classification. Among the sufficiently expressive hypothesis functions $h:\mathcal{X}\rightarrow[0,1]$ in a class $\mathcal{H}$, we define $h_{0}(\cdot)$ as a desired calibrated predictor for $y$, which minimizes the calibration error $\mathbb{E}_{x\sim\mathcal{D}}[(h(x)-c(x))^{2}]$ [39], where $c(x)=\mathbb{E}_{y}[y|h(x)]$ is the expected value of $y$ given a prediction $h(x)$. That is, $h_{0}$ always produces the calibrated prediction for $y$ given $x$ so that the output confidence $h_{0}(x)$ matches the expectation of $y$ over the subset of samples that have the same confidence value with $h_{0}(x)$. Our goal is to learn a hypothesis function $h(\cdot)$ that outputs reliable prediction probability on samples from the unseen OOD domain, which is defined by a distribution $\mathcal{D}_{\text{OOD}}$, as well as on the ID domain $\mathcal{D}_{\text{ID}}$, where the predictor is trained on. In essence, the error $\varepsilon_{\mathcal{D}}(h)=\mathbb{E}_{x\sim\mathcal{D}}[(h(x)-h_{0}(x))^{2}]$ should be small for two different domains $\mathcal{D}\in\{\mathcal{D}_{\text{ID}},\mathcal{D}_{\text{OOD}}\}$. Here, we focus on the covariate shift scenario (§2) that only the marginal distribution over $\mathcal{X}$ changes while the distribution over $\mathcal{Y}$ is preserved.
Let the optimal hypothesis $h^{*}$, which minimizes a combination of errors on both ID and OOD, be $h^{*}:=\operatorname*{arg\,min}_{h\in\mathcal{H}}\varepsilon_{\mathcal{D}_{\text{ID}}}(h)+\varepsilon_{\mathcal{D}_{\text{OOD}}}(h)$, and $\Delta$ denote the optimal joint error $\Delta:=\varepsilon_{\mathcal{D}_{\text{ID}}}(h^{*})+\varepsilon_{\mathcal{D}_{\text{OOD}}}(h^{*})$. Now, we derive a new bound for the OOD calibration error and OOD classification error.
@@
Let $h:\mathcal{X}\rightarrow[0,1]$ be a real-valued function of structure $h(x)=\sum_{i=1}^{d}h_{i}(x[i])$ where $h_{i}$ is an arbitrary one-dimensional function, and $h$ is in a hypothesis class $\mathcal{H}$ that has pseudo dimension $\mathcal{P}dim(\mathcal{H})=d_{h}$, $\hat{\mathcal{D}}_{\text{ID}}$ be an $N$-size empirical distribution on ID domain. If $(x[1],...,x[d])$ have matching marginals for ID and OOD, and $(x[i],x[j])$ is a bi-variate Gaussian for every $i,j\in[d]$, then for any $\delta\in(0,1)$ and for all $h$, the following bounds hold with probability at least $1-\delta$:
@@
where $\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}}{:=}\mathbb{E}_{\mathcal{D}_{\text{ID}}}[\tilde{x}\tilde{x}^{T}]$ is a covariance matrix with a strictly positive minimum singular value of $d$-dimensional normalized input $\tilde{x}=(\tilde{x}[1],...,\tilde{x}[d])$, where $\tilde{x}[i]{:=}(x[i]-\mathbb{E}[x[i]])\text{Var}(x[i])^{-1/2}$ and $\sigma_{\text{min}}(M)$ is the smallest singular value of a matrix $M\in\mathbb{R}^{d_{1}\times d_{2}}$. These theoretical results can be directly applied to the intermediate or penultimate layer’s representation of a neural network by setting the input variable $x$ as a representation vector as in [4, 73]. From now on, we will assume the input as an image representation from the last layer of the visual encoder in the following sections. Note that 1) the LHS of the first inequality (ineq.(11)) indicates OOD calibration error, 2) two terms in the LHS of the second inequality (ineq.(12)) denote OOD classification error in terms of $L_{2}$ loss and prediction sharpness on OOD domain, and 3) both inequalities have the same RHS, which contains the empirical estimate of ID calibration error, the reciprocal of the smallest singular value of ID input covariance matrix, and remaining irreducible terms that depend on the problem setup. Intuitively, Theorem 3.1 implies that pursuing diverse input features while maintaining the calibration of the classifier contributes to improving OOD calibration and generalization simultaneously. We defer the proof and discussion on the tightness of the bound and its assumptions in Appendix C.
@@ 4 Method
Our goal is to achieve good OOD generalization and calibration during the ID adaptation of pre-trained models. Motivated by Theorem 3.1, we propose a new fine-tuning method that increases the smallest singular value of the ID input covariance matrix while improving ID calibration, thereby lowering the upper bound of OOD calibration and generalization errors. By following [61, 30, 17], we set a vision-language model (VLM), CLIP [47] as our target, which serves as a remarkably strong backbone for zero-shot inference and fine-tuning with ease. We limit the scope of validation to image classification tasks. Note that our theorem is not confined to specific domains or model architectures and thus can be applied beyond CLIP’s image classification. See Figure 2 for an overview.